OpenStudy (anonymous):

Which of the following are solutions to the equation below? Check all that apply. 2x2 + 5x + 8 = 5 A. -1/2 B. 1 C. 3 D. -3 E. -1 F. -3/2

OpenStudy (anonymous):

start by subtracting 5 from both sides to get \[2x^2+5x+3=0\]

OpenStudy (anonymous):

then factor is \((x+1)(2x+3)=0\) and solve that one

OpenStudy (anonymous):

How?

OpenStudy (blaque678):

a/e

OpenStudy (anonymous):

\[x+1=0\iff x=-1\] \[2x+3=0\iff x=-\frac{3}{2}\]

OpenStudy (anonymous):

Okay. I get it know. Thanks